# Dynamic linear regression models for forecasting time series with semi long memory errors

### v1

@@ 3.1. Frequency domain likelihood
The frequency domain approach to inference relies on the asymptotic properties of the frequency representation of the time series process in (2.1). Engle, (1974) shows how to rewrite (2.1) in terms of the periodograms of $Y_{t}$ and $\boldsymbol{X}_{t}$ and derive the ordinary least squares estimator of the transformed regression, which is also shown to be the best linear unbiased estimator. Our Bayesian approach requires a log-likelihood function, which can be derived using asymptotic properties of the periodogram data, which we now outline in some detail.
@@
$Z_{t}=Y_{t}-\boldsymbol{X}_{t}^{\top}\boldsymbol{\beta},\,\,t=1,\dots,T.$
The general case assumes that $\eta_{t}$ in (2.1) is the seasonal ARTFIMA process in (2.8) and hence the spectral density is given by (2.9). The spectral density of the ARTFIMA process in (2.7) is obtained using $Q=P=0$ in (2.9). The spectral density of the seasonal ARMA process is obtained by setting $\lambda=0$ and $d=0$ in (2.9), and for ARMA, in addition, $Q=P=0$.
@@
$\mathcal{K}=\begin{cases}-\frac{T}{2},-\frac{T}{2}+1,\dots,\frac{T}{2}-1,\text{ if }T\text{ is even},\\ -\frac{(T-1)}{2},-\frac{(T-1)}{2}+1,\dots,\frac{(T-1)}{2},\text{ if }T\text{ is odd}.\end{cases}$
@@
(3.1) $J_{Z}(\omega_{k})$ $=\sum_{t=1}^{T}Z_{t}\exp(-i\omega_{k}t),$
which can be computed for all $T$ frequencies $\omega_{k}$, $k\in\mathcal{K}$, using $\mathcal{O}(T\log_{2}(T))$ operations via the fast Fourier transform (Cooley and Tukey,, 1965). Since the pseudo data depends on $\boldsymbol{\beta}$, the DFT in (3.1) needs to recomputed for each new sample in the Markov chain Monte Carlo (MCMC) algorithm. However, following Matsuda and Yajima, (2009),
$J_{Z}(\omega_{k})$ $=\sum_{t=1}^{T}(Y_{t}-\boldsymbol{X}_{t}^{\top}\boldsymbol{\beta})\exp(-i\omega_{k}t)$
$=\sum_{t=1}^{T}Y_{t}\exp(-i\omega_{k}t)-\left(\sum_{t=1}^{T}\boldsymbol{X}_{t}\exp(-i\omega_{k}t)\right)^{\top}\boldsymbol{\beta}$
@@
The DFT $J_{Z}(\omega_{k})$ is a weighted complex valued sum of the pseudo data. Peligrad and Wu, (2010, Theorem 2.1) show that under quite regular conditions,
(3.2) $\frac{1}{\sqrt{T}}(\Re(J_{Z}(\omega_{k})),\Im(J_{Z}(\omega_{k}))),\quad T\rightarrow\infty,$
where, respectively, $\Re(z)$ and $\Im(z)$ denote the real and imaginary parts of $z$, converge in distribution to a bivariate normal distribution with (asymptotically) independent components having expected value 0 and variance $\pi f(\omega_{k})$, with $f$ being the spectral density of $Z_{t}$. Moreover, they show that $\frac{1}{\sqrt{T}}J_{Z}(\omega_{k})$ are asymptotically independent for all $k\in\mathcal{K}$.
Define the periodogram
(3.3) $I_{Z}(\omega_{k})=\frac{1}{2\pi}\left|\frac{1}{\sqrt{T}}J_{Z}(\omega_{k})\right|^{2}.$
Then, for $k\in\mathcal{K}\setminus 0$, by (3.2),
(3.4) $\frac{I_{Z}(\omega_{k})}{f(\omega_{k})}$ $=\frac{1}{2}\left|\frac{1}{\sqrt{\pi f(\omega_{k})}}\frac{1}{\sqrt{T}}J_{Z}(\omega_{k})\right|^{2}\sim\frac{\chi^{2}(2)}{2},$
(3.5) $I_{Z}(\omega_{k})$ $\sim\mathrm{Exp}(f(\omega_{k})),\,\,k\in\mathcal{K}\setminus 0,$
where $\mathrm{Exp}$ denotes an exponential random variable parameterised by its mean. Emphasising the dependence on the parameters, the log-density of (3.5) is
(3.6) $\log p(I_{Z}(\omega_{k};\boldsymbol{\beta})|\boldsymbol{\theta})$ $=-\log(f(\omega_{k};\boldsymbol{\vartheta}))-\frac{I_{Z}(\omega_{k};\boldsymbol{\beta})}{f(\omega_{k};\boldsymbol{\vartheta})},\,\,k\in\mathcal{K}\setminus 0,$
where $\boldsymbol{\theta}=(\boldsymbol{\vartheta},\boldsymbol{\beta})$ contains all unknown parameters. When $k=0$, $J_{Z}(0)=0$ since $J_{Y}(0)=\sum_{t=1}^{T}Y_{t}=0$ and $\boldsymbol{J}_{\boldsymbol{X}}(0)=\sum_{t=1}^{T}\boldsymbol{X}_{t}=\boldsymbol{0}$ for demeaned data.
The asymptotic distributions of the periodogram data underlie the idea of the so-called Whittle log-likelihood (Whittle,, 1953): Use the log-density of each periodogram observation in $I_{Z}(\omega_{k};\boldsymbol{\beta})$ (3.6) to form the log-likelihood of all periodograms, which becomes a sum due to the (asymptotic) independence. For a real-valued process $Z_{t}$, both the periodogram and spectral density are symmetric around the origin; hence, only non-negative frequencies are considered. The Whittle log-likelihood is obtained by adding (due to the asymptotic independence) the log-densities in (3.6) for all the positive frequencies (zeroth frequency excluded with demeaned data). The Whittle log-likelihood is, for odd $T$,
(3.7) $\ell_{W}(\boldsymbol{\theta})$ $=-\sum_{k=1}^{(T-1)/2}\left(\log(f(\omega_{k};\boldsymbol{\vartheta}))+\frac{I_{Z}(\omega_{k};\boldsymbol{\beta})}{f(\omega_{k};\boldsymbol{\vartheta})}\right),$
and the summation runs to $T/2-1$ instead if $T$ is even. It can be shown that the Whittle log-likelihood is an approximation of the Gaussian time domain log-likelihood (Guyon,, 1982; Kent and Mardia,, 1996).
Finally, we stress that a frequency domain approach for dynamic linear regression models with the Whittle log-likelihood is not appropriate when using the ARFIMA model because the Whittle approximation fails to hold for long memory processes, in particular for the small frequencies (Robinson,, 1995; Rousseau et al.,, 2012). Thus, to carry out inference in dynamic linear regression models with ARFIMA errors, one has to resort to the time domain likelihood, which is considerably slower to compute compared to a frequency domain approach that uses the ARTFIMA process; see Sections 4.1, 5.4, and 5.5.
3.2. Bayesian inference via Markov chain Monte Carlo
An important objective in time series is to learn $\boldsymbol{\theta}=(\boldsymbol{\vartheta},\boldsymbol{\beta})\subseteq\boldsymbol{\Theta}\subset\mathbb{R}^{\dim_{\boldsymbol{\vartheta}}+\dim_{\boldsymbol{\beta}}}$ given realisations of the time series processes $Y_{t}\subseteq\mathbb{R}$ and $\boldsymbol{X}_{t}\subseteq\mathbb{R}^{\dim_{\boldsymbol{X}}}$. Let $p(\boldsymbol{Z}|\boldsymbol{\theta})$ denote the likelihood function of $\boldsymbol{\theta}$ given the pseudo data $\boldsymbol{Z}=(Z_{1},\dots,Z_{T})$, which depends on the subset $\boldsymbol{\beta}$ of the parameter vector $\boldsymbol{\theta}$, but we suppress this dependence for notational simplicity. When the likelihood is obtained via the Whittle log-likelihood in (3.7),  $p(\boldsymbol{Z}|\boldsymbol{\theta})=\exp\left(\ell_{W}(\boldsymbol{\theta})\right)$. The likelihood function in this case is given the periodogram ordinates $I(\omega_{0}),\dots,I(\omega_{(T-1)/2})$, but since they are functions of $\boldsymbol{Z}$ we keep using the notation $p(\boldsymbol{Z}|\boldsymbol{\theta})$.
The cornerstone of Bayesian inference is the posterior distribution obtained via Bayes’ theorem,
[…]
@@ 3.2. Bayesian inference via Markov chain Monte Carlo
where $p(\boldsymbol{\theta})$ is the prior distribution. MCMC methods are a class of iterative procedures to sample from (3.8), with the Metropolis-Hastings algorithm (Metropolis et al.,, 1953; Hastings,, 1970) arguably being the most used. The Metropolis-Hastings algorithm constructs a Markov chain $\left\{\boldsymbol{\theta}^{(j)}\right\}$ by starting at some initial value $\boldsymbol{\theta}^{(0)}$ and then, recursively, proposes a candidate draw $\boldsymbol{\theta}^{\prime}$ from a proposal density $q(\boldsymbol{\theta}|\boldsymbol{\theta}^{(j-1)})$ and sets $\boldsymbol{\theta}^{(j)}=\boldsymbol{\theta}^{\prime}$ with acceptance probability
@@
(3.9) $\alpha_{\mathrm{MH}}=\min\left(1,\frac{p(\boldsymbol{Z}|\boldsymbol{\theta}^{\prime})p(\boldsymbol{\theta}^{\prime})/q(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{(j-1)})}{p(\boldsymbol{Z}|\boldsymbol{\theta}^{(j-1)})p(\boldsymbol{\theta}^{(j-1)})/q(\boldsymbol{\theta}^{(j-1)}|\boldsymbol{\theta}^{\prime})}\right).$
If a proposed draw is rejected, $\boldsymbol{\theta}^{(j)}=\boldsymbol{\theta}^{(j-1)}$. The acceptance probability in (3.9) is set so that, informally speaking, the Markov chain consists of samples from the posterior in (3.8) after a warm-up period referred to as the burn-in. Let $\left\{\boldsymbol{\theta}^{(j)}\right\}_{j=1,\dots,N}$ be the samples after discarding the burn-in. The proposal density is often taken to be a random walk, e.g.
$q(\boldsymbol{\theta}|\boldsymbol{\theta}^{(j-1)})=\mathcal{N}(\boldsymbol{\theta}|\boldsymbol{\theta}^{(j-1)},c\boldsymbol{\Sigma}_{\mathrm{prop}}),$
where $\mathcal{N}(\boldsymbol{x}|\boldsymbol{\mu},\boldsymbol{\Sigma})$ denotes the multivariate normal density of $\boldsymbol{x}$ with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. Moreover, $\boldsymbol{\Sigma}_{\mathrm{prop}}$ is an approximation of the posterior covariance matrix and $c=2.38/\sqrt{\dim(\boldsymbol{\theta})}$ for optimality (Gelman et al.,, 1997).
@@
If the algorithm rejects too often, the Markov chain becomes “sticky”, which causes inefficient estimates of posterior expectations. To explain how this inefficiency is measured, suppose $\theta$ is scalar-valued and consider estimating the expectation of a function $h$, i.e.
@@
If the samples are independent, then the asymptotic variance of $\sqrt{N}\widehat{I}_{N}$ is $\sigma_{h}^{2}$, where $\sigma^{2}_{h}=\mathrm{Var}(h(\theta))$. However, MCMC results in correlated samples and then the asymptotic variance of $\sqrt{N}\widehat{I}_{N}$ is $\sigma_{h}^{2}(1+\sum_{i=1}^{\infty}\rho_{i})$, where $\rho_{i}$ is the autocorrelation at the $i$th lag the Markov chain. The term $(1+\sum_{i=1}^{\infty}\rho_{i})$ is called the inefficiency factor. It measures how much the asymptotic variance of the estimate of a posterior expectation is inflated compared to an ideal sampler that produces independent draws. The effective sample size (ESS) of the resulting Markov chain is defined as $\text{ESS}=N/(1+\sum_{i=1}^{\infty}\rho_{i})$. Thus, the effective sample size is an estimate of the number of samples equivalent to independent samples the sampler is generating.

### v2

@@ 3.1. Frequency domain likelihood
The frequency domain approach to inference relies on the asymptotic properties of the frequency representation of the time series process in (2.1). Engle, (1974) shows how to rewrite (2.1) in terms of the periodograms of $Y_{t}$ and $\boldsymbol{X}_{t}$ and derive the ordinary least squares estimator of the transformed regression, which is also shown to be the best linear unbiased estimator. Our Bayesian approach requires a log-likelihood function, which can be derived using asymptotic properties of the periodogram data, and is outlined below.
@@
$Z_{t}=Y_{t}-\boldsymbol{X}_{t}^{\top}\boldsymbol{\beta},\,\,t=1,\dots,T.$ (3.1)
The non-seasonal ARTFIMA process has the spectral density in (2.7), and when $\lambda=0$ and $d=0$, it simplifies to the spectral density for a non-seasonal ARMA model. The general case assumes that $\eta_{t}$ is a seasonal ARTFIMA process and has the spectral density given by (2.9). Similarly, the spectral density of the seasonal ARMA process is obtained by setting $\lambda=0$ and $d=0$.
@@
$\mathcal{K}=\begin{cases}-\frac{T}{2},-\frac{T}{2}+1,\dots,\frac{T}{2}-1,\text{ if }T\text{ is even},\\ -\frac{(T-1)}{2},-\frac{(T-1)}{2}+1,\dots,\frac{(T-1)}{2},\text{ if }T\text{ is odd}.\end{cases}$ (3.2)
@@
$J_{Z}(\omega_{k})$ $=\sum_{t=1}^{T}Z_{t}\exp(-\mathrm{i}\omega_{k}t),$ (3.3)
and can be computed for all $T$ frequencies $\omega_{k}$, $k\in\mathcal{K}$, using $\mathcal{O}(T\log_{2}(T))$ operations via the fast Fourier transform (Cooley and Tukey,, 1965). Since the pseudo data depends on $\boldsymbol{\beta}$, the DFT in (3.3) needs to be recomputed for each new sample in the Markov chain Monte Carlo (MCMC) algorithm. However, following Matsuda and Yajima, (2009),
$J_{Z}(\omega_{k})$ $=\sum_{t=1}^{T}(Y_{t}-\boldsymbol{X}_{t}^{\top}\boldsymbol{\beta})\exp(-\mathrm{i}\omega_{k}t)$
$=\sum_{t=1}^{T}Y_{t}\exp(-\mathrm{i}\omega_{k}t)-\left(\sum_{t=1}^{T}\boldsymbol{X}_{t}\exp(-\mathrm{i}\omega_{k}t)\right)^{\top}\boldsymbol{\beta}$
@@
where $J_{Y}$ is the DFT of $Y_{t}$, and $\boldsymbol{J}_{\boldsymbol{X}}(\omega_{k})$ is the $m$-dimensional column vector with the $j$th element being the DFT of the (univariate) time series $X_{jt}$. Thus, since $J_{Y}(\omega_{k})$ and $\boldsymbol{J}_{\boldsymbol{X}}(\omega_{k})$ only depend on the data $Y_{t}$ and $\boldsymbol{X}_{t}$, and can be pre-computed before the MCMC algorithm, $J_{Z}(\omega_{k})$ can be evaluated in $\mathcal{O}(T)$ for all frequencies when $\boldsymbol{\beta}$ changes (assuming $m\ll T$).
Let $\Re(J_{Z}(\omega_{k}))$ and $\Im(J_{Z}(\omega_{k}))$ be the real and imaginary parts of the DFT $J_{Z}(\omega_{k})$ which is a weighted complex valued sum of the pseudo data. Peligrad and Wu, (2010, Theorem 2.1) show that under some regularity conditions,
$\frac{1}{\sqrt{T}}(\Re(J_{Z}(\omega_{k})),\Im(J_{Z}(\omega_{k}))),\quad T\rightarrow\infty,$ (3.4)
converge in distribution to a bivariate normal distribution with (asymptotically) independent components having expected value 0 and variance $\pi f(\omega_{k})$, with $f$ being the spectral density of $Z_{t}$. Moreover, they show that the $\frac{1}{\sqrt{T}}J_{Z}(\omega_{k})$ are asymptotically independent for all $k\in\mathcal{K}$.
Define the periodogram
$I_{Z}(\omega_{k})=\frac{1}{2\pi}\left|\frac{1}{\sqrt{T}}J_{Z}(\omega_{k})\right|^{2}.$ (3.5)
Then, for $k\in\mathcal{K}\setminus 0$, by (3.4),
$\frac{I_{Z}(\omega_{k})}{f(\omega_{k})}$ $=\frac{1}{2}\left|\frac{1}{\sqrt{\pi f(\omega_{k})}}\frac{1}{\sqrt{T}}J_{Z}(\omega_{k})\right|^{2}\sim\frac{\chi^{2}(2)}{2},$ (3.6)
$I_{Z}(\omega_{k})$ $\sim\mathrm{Exp}(f(\omega_{k})),\,\,k\in\mathcal{K}\setminus 0,$ (3.7)
where $\mathrm{Exp}(f(\omega_{k}))$ denotes an exponential distribution with mean $f(\omega_{k})$. Emphasising the dependence on the parameters, the log-density of (3.7) is
$\log p(I_{Z}(\omega_{k};\boldsymbol{\beta})|\boldsymbol{\theta})$ $=-\log(f(\omega_{k};\boldsymbol{\vartheta}))-\frac{I_{Z}(\omega_{k};\boldsymbol{\beta})}{f(\omega_{k};\boldsymbol{\vartheta})},\,\,k\in\mathcal{K}\setminus 0,$ (3.8)
where $\boldsymbol{\theta}=(\boldsymbol{\vartheta}^{\top},\boldsymbol{\beta}^{\top})^{\top}$ contains all unknown parameters. When $k=0$, $J_{Z}(0)=0$ since $J_{Y}(0)=\sum_{t=1}^{T}Y_{t}=0$ and $\boldsymbol{J}_{\boldsymbol{X}}(0)=\sum_{t=1}^{T}\boldsymbol{X}_{t}=\boldsymbol{0}$ for demeaned data.
The asymptotic distributions of the periodogram data underlie the idea of the so-called Whittle log-likelihood (Whittle,, 1953): use the log-density of each periodogram observation $I_{Z}(\omega_{k};\boldsymbol{\beta})$ in (3.8) to form the log-likelihood of all periodograms, which becomes a sum due to the (asymptotic) independence. For a real-valued process $Z_{t}$, both the periodogram and spectral density are symmetric around the origin; hence, only non-negative frequencies are considered. The Whittle log-likelihood is obtained by adding (due to the asymptotic independence) the log-densities in (3.8) for all the positive frequencies (zeroth frequency excluded with demeaned data). The Whittle log-likelihood is, for odd $T$,
$\ell_{W}(\boldsymbol{\theta})$ $=-\sum_{k=1}^{(T-1)/2}\left(\log(f(\omega_{k};\boldsymbol{\vartheta}))+\frac{I_{Z}(\omega_{k};\boldsymbol{\beta})}{f(\omega_{k};\boldsymbol{\vartheta})}\right),$ (3.9)
and the summation runs to $T/2-1$ instead if $T$ is even. Guyon, (1982) and Kent and Mardia, (1996) show that the Whittle log-likelihood is a spectral approximation of the Gaussian time domain log-likelihood. In particular, they derive probabilistic convergence results of, respectively, the Whittle log-likelihood and its derivative, to the Gaussian log-likelihood and its derivative.
Finally, we stress that a frequency domain approach via the Whittle log-likelihood for dynamic linear regression models is inappropriate when using the ARFIMA model, because the Whittle approximation fails to hold for long memory processes, in particular for the small frequencies (Robinson,, 1995; Rousseau et al.,, 2012). Thus, to carry out inference in dynamic linear regression models with ARFIMA errors, it is necessary to use the time domain log-likelihood, which is considerably slower to compute compared to a frequency domain approach that uses the ARTFIMA process; see Sections 4.1, 5.4, and 5.5.
3.2. Bayesian inference via Markov chain Monte Carlo
An important ain in time series is to learn the posterior distribution of $\boldsymbol{\theta}=(\boldsymbol{\vartheta}^{\top},\boldsymbol{\beta}^{\top})^{\top}\subseteq\boldsymbol{\Theta}\subset\mathbb{R}^{\dim_{\boldsymbol{\vartheta}}+\dim_{\boldsymbol{\beta}}}$ given realisations of the time series processes $Y_{t}\in\mathbb{R}$ and $\boldsymbol{X}_{t}\in\mathbb{R}^{\dim_{\boldsymbol{X}}}$. Let $p(\boldsymbol{Z}|\boldsymbol{\theta})$ denote the likelihood function of $\boldsymbol{\theta}$ given the pseudo data $\boldsymbol{Z}=(Z_{1},\dots,Z_{T})^{\top}$, which depends on the subset $\boldsymbol{\beta}$ of the parameter vector $\boldsymbol{\theta}$, but we suppress this dependence for notational simplicity. When the likelihood is obtained via the Whittle log-likelihood in (3.9),  $p(\boldsymbol{Z}|\boldsymbol{\theta})=\exp\left(\ell_{W}(\boldsymbol{\theta})\right)$. The likelihood function in this case is given the periodogram ordinates data $\boldsymbol{I}=(I(\omega_{0}),\dots,I(\omega_{(T-1)/2}))^{\top}$, but since they are functions of $\boldsymbol{Z}$ we keep using the notation $p(\boldsymbol{Z}|\boldsymbol{\theta})$ (rather than $p(\boldsymbol{I}|\boldsymbol{\theta})$).
The posterior distribution of $\boldsymbol{\theta}$ is obtained using Bayes’ theorem,
@@ 3.2. Bayesian inference via Markov chain Monte Carlo
where $p(\boldsymbol{\theta})$ is the prior distribution of $\boldsymbol{\theta}$. Markov chain Monte Carlo methods are a class of iterative procedures to sample from (3.10), with the Metropolis-Hastings algorithm (Metropolis et al.,, 1953; Hastings,, 1970) arguably being the most popular. The Metropolis-Hastings algorithm constructs a Markov chain $\left\{\boldsymbol{\theta}^{(j)}\right\}$ by starting at some initial value $\boldsymbol{\theta}^{(0)}$ and then, recursively, proposes a candidate draw $\boldsymbol{\theta}^{\prime}$ from a proposal density $q(\boldsymbol{\theta}|\boldsymbol{\theta}^{(j-1)})$ and sets $\boldsymbol{\theta}^{(j)}=\boldsymbol{\theta}^{\prime}$ with acceptance probability
@@
$\alpha_{\mathrm{MH}}=\min\left(1,\frac{p(\boldsymbol{Z}|\boldsymbol{\theta}^{\prime})p(\boldsymbol{\theta}^{\prime})/q(\boldsymbol{\theta}^{\prime}|\boldsymbol{\theta}^{(j-1)})}{p(\boldsymbol{Z}|\boldsymbol{\theta}^{(j-1)})p(\boldsymbol{\theta}^{(j-1)})/q(\boldsymbol{\theta}^{(j-1)}|\boldsymbol{\theta}^{\prime})}\right).$ (3.11)
If a proposed draw is rejected, $\boldsymbol{\theta}^{(j)}=\boldsymbol{\theta}^{(j-1)}$. The proposal density is often a random walk, e.g.
$q(\boldsymbol{\theta}|\boldsymbol{\theta}^{(j-1)})$ $=\mathcal{N}(\boldsymbol{\theta}|\boldsymbol{\theta}^{(j-1)},c\boldsymbol{\Sigma}_{\mathrm{prop}}),$ (3.12)
where $\mathcal{N}(\boldsymbol{x}|\boldsymbol{\mu},\boldsymbol{\Sigma})$ denotes the multivariate normal density of $\boldsymbol{x}$ with mean $\boldsymbol{\mu}$ and covariance matrix $\boldsymbol{\Sigma}$. Moreover, $\boldsymbol{\Sigma}_{\mathrm{prop}}$ is an approximation of the posterior covariance matrix, for example minus the inverse of the observed Fisher information evaluated at the maximum a posteriori (MAP) estimate, and $c=2.38/\sqrt{\dim(\boldsymbol{\theta})}$ for optimality, resulting in $\alpha_{\mathrm{MH}}=0.234$ (Gelman et al.,, 1997). However, in the presence of intricate geometry in the posterior due to complex dependence structures, selecting a proposal covariance as described may yield low acceptance probabilities. An alternative approach that we implement is to adaptively tune the proposal covariance matrix (Haario et al.,, 2001). This technique uses the information accumulated thus far through the previous draws to adjust the covariance proposal matrix. The adaptive approach can lead to more accurate simulated approximations of the target posterior and less auto-correlation in the resulting chain. To target a specified overall acceptance rate of the sampler (0.234), we follow Garthwaite et al., (2016) who utilise a Robbins-Monro search process to compute the scaling constant $c$ in (3.12) of the adapted proposal covariance matrix.
@@
The acceptance probability in (3.11) is set so that, informally speaking, the Markov chain consists of samples from the posterior in (3.10) after a warm-up period referred to as the burn-in. Let $\left\{\boldsymbol{\theta}^{(j)}\right\}_{j=1,\dots,N}$ be the samples after discarding the burn-in. If the algorithm rejects too often, the Markov chain (the samples) becomes “sticky”, which causes inefficient estimates of posterior expectations. To explain how this inefficiency is measured, suppose $\theta$ is scalar-valued and consider estimating the expectation of a function $h$, i.e.
@@
If the samples are independent, then the asymptotic variance of $\sqrt{N}\widehat{\mathcal{I}}_{N}$ is $\sigma_{h}^{2}$, where $\sigma^{2}_{h}=\mathrm{Var}(h(\theta))$. However, MCMC results in correlated samples and then the asymptotic variance of $\sqrt{N}\widehat{\mathcal{I}}_{N}$ is $\sigma_{h}^{2}(1+\sum_{i=1}^{\infty}\rho_{i})$, where $\rho_{i}$ is the autocorrelation at the $i$th lag the Markov chain. The term $(1+\sum_{i=1}^{\infty}\rho_{i})$ is called the inefficiency factor. It measures how much the asymptotic variance of the estimate of a posterior expectation is inflated compared to an ideal sampler that produces independent draws. The effective sample size (ESS) of the resulting Markov chain is defined as $\text{ESS}=N/(1+\sum_{i=1}^{\infty}\rho_{i})$ (note that $\text{ESS}=N$ with independent samples). Thus, the effective sample size estimates the number of samples equivalent to the number of independent samples the sampler is generating, see Geyer, (2011) for more details.